### v1

# Received August 1, 2010; revised October 1, 2010; accepted for publication November 1, 2010

Abstract
Uncertainty around predictions from a model due to the finite size of the development sample has traditionally been approached using classical inferential techniques. The finite size of the sample will result in the discrepancy between the final model and the correct model that maps covariates to predicted risks. From a decision-theoretic perspective, this discrepancy might affect the subsequent treatment decisions, and thus is associated with utility loss. From this perspective, procuring more development data is associated in expected gain in the utility of using the model. In this work, we define the Expected Value of Sample Information (EVSI) as the expected gain in clinical utility, defined in net benefit (NB) terms in net true positive units, by procuring a further development sample of a given size. We propose a bootstrap-based algorithm for EVSI computations, and show its feasibility and face validity in a case study. Decision-theoretic metrics can complement classical inferential methods when designing studies that are aimed at developing risk prediction models. predictive analytics; precision medicine; decision theory; value of information; bayesian statistics
00footnotetext: Correspondence to Abdollah Safari, University of Tehran, Tehran, Iran; email: a.safari@ut.ac.ir
1 Introduction
Risk prediction models that utilize patient characteristics to predict the risk of clinically relevant events are important enablers of precision medicine and individualized care (Kattan and others, 2016; Steyerberg, 2019; Van Calster and others, 2013). Typically, such models are classical regression-based or machine learning algorithms developed using a finite ’development’ sample from a relevant target population. Examples of such models are the Framingham Risk Score for predicting the risk of cardiovascular outcomes (D’Agostino Sr and others, 2008), or the Adnex score for predicting the risk of malignancy in ovarian tumors (Timmerman and others, 2016).
Predictions from such models are inherently subject to uncertainty, primarily due to the finite size of the development sample (Harrell, 2015; Steyerberg, 2019). Conventionally, such uncertainty is communicated through standard frequentist methods around performance metrics of the model (e.g., 95% confidence interval around c-statistic or calibration line [DeLong and others, 1988]). Recent developments have focused on studying the stability of predicted risks, with proposals on the graphical and statistical presentations of variation in predicted risks (Riley and Collins, 2023). Similarly, in designing risk prediction development studies, contemporary approaches for sample size calculations are based on statistical inference on classical metrics of model performance. For example, Riley and others (2024) have proposed multi-criteria sample size rules targeting prediction error, calibration, and discrimination. These significant developments have provided much needed objectivity in the design of risk prediction development and validation studies.
In addition to classical metrics of model performance around calibration and discrimination, the clinical utility of a model, expressed in terms of net benefit (NB), has become a standard metric of model performance (Vickers and Elkin 2006a; Efthimiou and others 2024). Among conventional metrics of model performance, net benefit (NB) is a decision-theoretic one, in the sense that the use of a risk prediction model is clinically justified if its expected NB is above the expected NB of the alternative decisions. Decision theory provides a fundamentally different approach towards uncertainty quantification than methods based on conventional statistical inference (Claxton, 1999). From the perspective of this theory, uncertainty results in the discrepancy between the predicted and correct (calibrated) risks, which in turn might result in an incorrect medical decision (a decision based on predicted risk that is different from when knowing the correct risk). Thus, from this perspective, uncertainty can be associated with a loss of clinical utility (Steyerberg, 2019). Such decision-theoretic approach towards evaluating the consequences of uncertainty in decisions, generally referred to as Value of Information (VoI) analysis, has been employed in various fields, such as decision-analytic modeling (Jackson and others, 2022; Jackson and others, 2019), environmental risk analysis (Yokota and Thompson, 2004; Weinstein, 1980; O’Hagan and others, 2006), and economic trials (Everest and others, 2023). Recently, such methods have been adapted to predictive analytics (Jackson and others, 2022; Vickers and others, 2023). Specifically, the expected value of perfect information (EVPI) has been applied to both development and validation phases of risk prediction models (Sadatsafavi and others, 2022; Sadatsafavi and others, 2023a). In the context of model development, EVPI quantifies the expected utility gain over a model developed using a finite sample by completely knowing the correct (strongly calibrated) model.
However, uncertainty can never be entirely eliminated. Procuring more samples for model development is expected to reduce, but not eliminate, prediction uncertainty (Greenland, 2000). As such, the EVPI puts an upper limit on the expected gain from future validation studies. In decision theory, the expected value of sample information (EVSI) quantifies the potential value of collecting additional data to inform a specific research design (Claxton, 1999; Willan and Briggs, 2006). In health policy-making, EVSI has been applied to quantify the expected utility (net benefit) gain from conducting a clinical trial of a given design and sample size (Ades and others, 2004a). Such expected gain can be juxtaposed with the costs of the trial to optimize the sample size (Willan and others, 2012).
The EVSI has recently been proposed for the validation phase of a model (evaluating the performance of a pre-specified model in a new population without changing its structure [Sadatsafavi and others, 2023b]). To the best of our knowledge, it has not yet been applied to the development phase of such models . Recent best practice standards recommend for not splitting the available data into development and validation, and instead use all the data for development [Collins and others, 2024]. As well, in many instances where a model is being validated in a new population, some form of model updating will be required (ranging from changing the intercept, to refitting the model in the new population [Binuya and others, 2022]). These mean in many practical instances, the investigators need to assess the NB of their model in the same data that is used to fit or update the model. As such, informing sample size calculations for model development around NB is a relevant pursuit.
The objective of this paper is to define EVSI for this context and propose algorithms for its computation. The rest of this paper is organized as follows. Section 2 contains context and notations, method and main findings of the paper. Section 3 includes the numerical results of our simulation study as well as evaluation of our method on a real dataset. Finally, Section 4 contains our conclusions and discussion.
2 Context and Notations
We focus on the development phase of a risk prediction model for a binary outcome (e.g., whether an individual with a heart attack will die in the next 30 days). We assume the investigator has had access to $n$ iid observations of predictors and response values from the relevant target population to develop a prediction model for the conditional risk of the outcome given a set of predictor values. Due to the finite size of the development sample, predictions from such a model might not be accurate even if the model form is correctly specified, resulting in potential loss of clinical utility. Procuring a further development sample can improve the accuracy of predictions, and reduce such a loss. The figure of merit is the expected gain in clinical utility from procuring a future development sample $D^{*}$ of size $n^{*}$ from the same target population, i.e., development EVSI. We assume the future sample will be exchangeable with the current sample (thus the investigator will naturally combine the current and future samples to refit their model). We quantify clinical utility in terms of net benefit (NB), as quantified in decision curve analysis (Vickers and Elkin, 2006b).
2.1 Notations
Let $D=(\mathbf{X},Y)=\{(\mathbf{X_{i}},Y_{i})\}_{i=1}^{n}$ be the current development sample where $\mathbf{X}$ represents the vector of predictors and $Y$ the binary outcome, and consequently $d=(\mathbf{x},y)=\{(\mathbf{x_{i}},y_{i})\}_{i=1}^{n}$ be its realization after observing the sample. The development data are presumed to obey the following joint distribution
$P_{\theta,\eta}(X,Y)=P_{\eta}(X)P_{\theta}(Y|X),$ (1)
where $\eta$ and $\theta$ are the parameters indexing, respectively, the marginal distribution of $X$ (i.e., the case-mix) and the conditional distribution of Y given X ($Y|X$, i.e., the response model). Given the binary nature of $Y$, let $\pi(x;\theta)=P(Y=1|X=x)$ be the risk of outcome for patient with characteristics $x$, and consequently, $\pi(X;\theta)=P(Y=1|X)$ be a random draw from the distribution of predicted risks specified by a model parameterized by $\theta$. Related, $\hat{\eta}$ and $\hat{\theta}$ estimate the case mix and model coefficients from the available data $d$, respectively. In what follows, we parametrically model $Y$, estimating $\theta$ (regression coefficients), in a general regression framework, while we adopt a non-parametric approach towards modeling case-mix, where $\hat{\eta}$ is the empirical distribution of $\mathbf{X}$.
2.2 Net benefit
We determine clinical utility by using NB, a widely recognized method for assessing the performance of risk prediction models due to its solid theoretical foundation and straightforward application Vickers and Elkin, 2006a. When converting continuous risk predictions into binary decisions, such as deciding whether or not to treat a patient, it is necessary to set a treatment threshold that aligns with the specific clinical context. This threshold should ideally reflect the relative importance of avoiding unnecessary treatments (false positives) versus ensuring necessary treatments (true positives). As explained by Vickers and Elkin, 2006a, this threshold effectively functions as a balancing point between true positives and false positives, allowing NB to be calculated by comparing the benefits of correct treatment decisions against the drawbacks of incorrect ones.
To compute NB, a particular risk threshold, denoted by $z$, is chosen. Patients with a predicted risk below $z$ are considered low-risk and are not treated, while those with a risk above $z$ are classified as high-risk and are treated. For those whose predicted risk is exactly $z$, there is no strong preference for treatment or non-treatment, as treating these patients would lead to a certain proportion being appropriately treated (true positives) and the rest being treated unnecessarily (false positives). This balance suggests that the negative impact of a false positive is weighted by $\frac{z}{1-z}$ in relation to a true positive. By using this threshold, the model’s true and false positive rates can be calculated, and NB can be expressed as the net gain in true positives after accounting for the cost of false positives. This approach provides a flexible way to evaluate NB across different thresholds $z$ and can be expressed as follows (dependency on $z$ omitted for simplicity)
NB of $\text{NB of model}=P(\text{True Positive})-P(\text{False Positive})\frac{z}{1-z}$ (2)
where, as mentioned earlier, $\frac{z}{1-z}$ represents the relative weight of a false-positive versus a true-positive classification and thus captures the harm-benefit trade-off at this threshold. In practice, the NB is often calculated for a plausible range of thresholds.
To make a decision whether the model is useful in this population, we shall compare its NB against the NB of at least two default decisions: treating all and treating none. The latter has 0 NB by definition, and the NB of treating all can be written as
NB of treating $\text{NB of treating all}=\text{Prevalence}-(1-\text{Prevalence})\frac{z}{1-z}.$ (3)
We index these three strategies as 0=treat none, 1=use model, and 2=treat all. Then the true NB function for a model indexed by $\hat{\theta}$ can generally be defined as (for brevity we drop the notation that indicates the NB quantities are dependent on $z$):
$NB\left(i;\hat{\theta},\eta,\theta\right)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \mathbb{E}\left\{I\left(\pi\left(X,\hat{\theta}\right)\geq z\right)\left[\pi(X,\theta)-(1-\pi(X,\theta))\frac{z}{1-z}\right]\right\}&i=1\text{ (use model)}\\ \mathbb{E}\{\pi(X,\theta)-(1-\pi(X,\theta))\frac{z}{1-z}\}&i=2\text{ (treat all)}\end{array}\right.,$ (4)
with $I(\cdot)$ being the indicator function.
The developments proposed here are also applicable to the situations where more than one model is being evaluated. We do not present this for the ease of presentation but the general principles will be the same.
In the current development sample $d$ of size $n$, given $\theta$ and $\hat{\theta}$ while plugging in the empirical estimate of $\eta$ (a weight of $1/n$ assigned to each observation), a consistent estimate of NBs can be written as:
$\hat{NB}\left(i;\hat{\theta},\theta\right)=\left\{\begin{array}[]{ll}0&i=0\\ \frac{1}{n}\sum_{j=1}^{n}\left(I\left(\pi\left(x_{j},\hat{\theta}\right)>z\right)\left[\pi(x_{j},\theta)-(1-\pi(x_{j},\theta))\frac{z}{1-z}\right]\right)&i=1\\ \frac{1}{n}\sum_{j=1}^{n}\left(\left[\pi(x_{j},\theta)-(1-\pi(x_{j},\theta))\frac{z}{1-z}\right]\right)&i=2\end{array}\right.$ (5)
where $\pi(x_{j},\theta)$ and $\pi\left(x_{j},\hat{\theta}\right)$ are the true and predicted risk of the $j^{th}$ patient, respectively.
2.3 Expected value of current information
The true NBs formulated in the previous section (equation 4) all depend on the true parameters $\eta$ and $\theta$, and their estimate expressed in 5, after substituting $\eta$ by its empirical distribution, still depends on $\theta$. However, we do not know the true value of $\eta$ and $\theta$, and instead know about their likely values via $P(\eta,\theta|d)$. One can compute the expected NB (ENB) of the three strategies with respect to this distribution as
$ENB(i;d)=\mathbf{E}_{\eta,\theta|d}\left[NB\left(i;\hat{\theta},\eta,\theta\right)\right]$ (6)
for $i=0,1,2$, where $E_{\eta,\theta|d}$ represents the expectation with respect to the posterior distribution of $P(\eta,\theta|d)$.
Consequently, decision making under current info can be formulated as follows: given sample $d$, a decision maker fits a model to obtain the model parameter estimate $\hat{\theta}$, estimates the ENB of the model, and declares a winner $W(d)$, which returns a value in $\{0,1,2\}$. The most rational approach for a risk-neutral decision maker is to choose the winning strategy as
$W(d)=\operatorname*{arg\,max}_{i}\{ENB(i;d)\}$ (7)
To compute $ENB(i;d)$ for model development, we propose a generic bootstrap-based algorithm as follows. First take a set of weights for the development sample $d$ as the Bayesian bootstrap sample $d^{(t)}=\left(x^{(t)},y^{(t)}\right)$, a weighted version of the original development sample. Next, refit the same original risk prediction model (a weighted model) on $d^{(t)}$ to obtain the model coefficient estimates $\theta^{(t)}$ and consequently, the vector of predicted risks $\pi^{(t)}=\pi\left(x^{(t)},\theta^{(t)}\right)$ from the fitted model. We treat $d^{(t)}$ as the population and $\theta^{(t)}$ as the vector of true model parameters, and $\pi^{(t)}$ as the vector of true risks. One can plug in these values in (5) to obtain the NB of different strategies on the development sample. To estimate ENBs and consequently compute EVSI, we simply need to repeat the above steps $T$ times, $t=1,\dots,T$, and then use the average of NBs as their estimated expectation, that is
$\hat{ENB}(i;d)=\left\{\begin{array}[]{ll}0&i=0\text{ (treat none)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(1;\hat{\theta},\theta^{(t)}\right)&i=1\text{ (use model)}\\ \frac{1}{T}\sum_{t=1}^{T}\hat{NB}\left(2;\hat{\theta},\theta^{(t)}\right)&i=2\text{ (treat all)}\end{array}\right.,$ (8)
2.4 Expected value of sample information (EVSI)
Assume for the moment that we know the true case-mix ($\eta$) and the correct model ($\theta$). As such, we can generate data of the future study $D^{*}$ (Heath and others, 2021). Given the assumption on the exchangability of $D$ and $D^{*}$, $D^{*}$ would be combined with the $D$, which we denote by $D^{+}=(D,D^{*})$, and the investigator would revisit their decision for declaring the superior decision based on $W(D^{+})$. The ENB associated with a given $D^{+}$ is $ENB\left(W(D^{+});D^{+}\right)$.
However, as the future data have not been collected, we will need to take expectation with respect to the predictive distribution of $D^{*}$:
$EVSI=$ $E_{D^{+}}\left[ENB\left(W(D^{+});D^{+}\right)\right]-ENB(W(D);D)$ (9)
EVSI is a scalar quantity in the same NB unit as the clinical utility of the model. The higher the EVSI, the higher the expected benefit from a sample. The EVSI asymptotes to EVPI.
2.5 Bootstrap-based EVSI computation
The second term on the EVSI formula in (9) is the ENB and can be estimated following the same generic bootstrap algorithm given in Section 2.3. Therefore, to compute EVSI, we need to estimate the first term. As the equation for EVSI implies, the nested computation of the two expectations generally gives rises to nested Monte Carlo algorithms (Ades and others, 2004b; Kunst and others, 2020). Our non-parametric specification of $\eta$ gives rise to a nested algorithm based on the Bayesian bootstrap. We note that, because the Bayesian bootstrap allows one to interpret $D^{t}$ as a random draw from the population that has generated $d$, we can use $D^{t}$ to also generate the sample for the planned future study. We can generate a random draw from $D^{*}$ through another level of sampling. Specifically, let $m$ be the size of the future sample. We then take a (second level of) bootstrap sample from $D^{(t)}$, called $D^{(*t)}=(X^{(*t)},Y^{(*t)})$, as our future sample of size $n^{*}$ given $\theta^{(t)}$ (and consequently $\pi^{(t)}$). Note that, since we drew $D^{(*t)}$ under the assumption of $\pi^{(t)}$s being the true risks for the population individuals, we generate $Y^{(*t)}_{i}$s from Bernoulli distributions with their corresponding true risks being the success probabilities. Next, we update the original prediction risk model based on the pooled dataset of $D^{(+t)}=(D,D^{(*t)})$ as well as the predicted risk vector of $\pi\left(X^{(+t)},\hat{\theta}^{(t)}\right)$ obtained from the updated model. The bootstrap-based EVSI computation algorithm presented in Algorithm 1.
Algorithm 1 EVSI & EVPI
  for $t=1,\dots,T$ do
     A random draw from the generating population $D^{(t)}$: bootstrap sample from original dataset
     True risks $\mathbf{\pi}^{(t)}$: estimated risks from the refitted model on $D^{(t)}$
     True NBs: estimate according to (5) on the development sample
     True NB of perfect information: estimate according to (5) where substitute $\pi\left(x_{j},\hat{(}\theta)\right)$ by its true value  [^1] See Sadatsafavi and others (2023b) for more details
     Future sample $D^{(*t)}$: bootstrap sample of size $m$ from $D^{(t)}$
     true outcomes: replace outcomes in $D^{(*t)}$ by a generated set of $Y^{(*t)}$ based on $\pi^{(t)}$
     Merged sample $D^{(+t)}$: merge the future sample $D^{(*t)}$ with the development sample $D$
     Updated risks $\pi\left(X^{(+t)},\hat{\theta}^{(t)}\right)$: estimated risks from the refitted model on $D^{(+t)}$
     Estimate NBs: estimate according to (5) on the merged sample
  end for
  Estimated ENBs: average estimated NBs based on $D$ according to (8)
  Estimated ENBs: average estimated NBs based on $D^{+}$s according to (8)
  Winner strategy: identify the strategy with maximum estimated ENB obtained from the previous step
  EVPI: Compute as the difference between ENB of perfect information and ENB of winner strategy
  EVSI: Compute according to (9)
3 Case study: GUSTO trial
As a case study, we apply our findings to GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI, or ’heart attack’) (investigators, 1993). The GUSTO trial was a four-arm randomized clinical trial that compared the efficacy of different intravenous thrombolytic regimens for AMI. This dataset has been widely used for methodological research in predictive analytics (Ennis and others, 1998). We use these data to build a prediction model for predicting 30-day mortality for AMI. We used sex, age, diabetes (yes or no), milocc (which part of the heart muscle is affected by the blocked artery and coded as 3=anterior, 4=inferior, 5=other, and 6=no MI), previous MI, hypertension, smoking, killip class, and treatment as covariates in the model. We consider risk threshold of 2% as the default threshold, but will evaluate the model at the entire possible range of thresholds.
Given the relatively large sample size (n=40830), the uncertainty in predictions when the model is fitted using the full sample is negligible. Indeed, a previous study reported an EVPI of 0 with the full dataset (thus, EVSI will be zero for all sample sizes [Sadatsafavi and others, 2023a]). In a similar fashion as Sadatsafavi and others (2023a), we took, without replacement, a random sub-sample of 1000 patients from GUSTO data and treat that as our development sample (current information).
Figure 1: Estimated NB of the model and the alternative default strategies (treat all and treat none) at different threshold values (panel A) and estimated NB increments of model versus the best alternative (panel B).
We start by drawing the decision curve at the entire possible range of thresholds, as shown in Figure 1 (panel A). The figure shows the NB of the model as well as the alternative default strategies (treat none and treat all). The model had higher expected NB than the alternative strategies at the all threshold values of interest. Additionally, panel (B) of Figure 1 shows the difference between the estimated NB of the model and that of the best alternative default strategy, i.e., $ENB(1,d)-max\left(0,ENB(2,d)\right)$, along with its bootstrap $95\%$ CI. Again, the model had a higher estimated NB that the default alternatives at the entire range of threshold values of interest.
Next, we draw EVSI. We chose four different future sample sizes: 100, 1000, 5000, and 10000. Following algorithm 1, for each sample size, we drew $10^{5}$ bootstrap samples ($T=10^{5}$) from the development sample, and EVSI was evaluated at every 0.01 point ($z$ values) between 0 and 1 (possible range for risks). Figure 2 illustrates computed EVSI for different sample sizes (coloured curves) at different threshold values, $z$, compared to the computed EVPI (black curve).
Figure 2: Computed EVSI (dotted/dashed curves) for future samples with four different seizes compared to the computed EVPI (solid curve).
EVPI quantifies the expected NB loss per decision (every time the model might be used). The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made at the target population (Willan and others, 2012). At the 0.21 threshold, a future validation study can have a maximum benefit equal to 0.003. One can put the VoI gain into context of the target population following the scheme proposed by Sadatsafavi and others (2023a). There are around 800,000 AMIs in the US every year (Virani and others, 2021). That is, at the 0.21 threshold value, a future validation study can have a maximum benefit equal to gaining 2050 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 7709 extra false positive cases per year (individuals who will not die within 30 days).
As expected, the larger the future samples, the greater EVSI, and consequently, the closer curves to that of EVPI. The overall pattern of computed EVSIs remain similar to that of EVPI for different sample sizes across threshold values. The ‘diminishing return’ pattern is also obvious for both thresholds: the expected NB gain steeply rises at small samples and plateaus as $n$ grows. At 0.21 threshold, a future study of size $n=1000$ has a per-decision EVSI value of 0.00132, corresponding to population value of 1058 in true positive cases gained (or 3978 in false positive cases averted), while a future study of size $n=100000$ has a per-decision EVSI value of 0.00234, corresponding to population value of 1869 in true positive cases gained (or 7030 in false positive cases averted).
4 Discussion
In this study, we applied the value of information (VoI) framework to assess the decision-theoretic consequences of uncertainty due to the finite size of the development sample. We introduced the expected value of sample information (EVSI) for model development, defining it as the anticipated increase in net benefit (NB) gained from obtaining an additional development sample of a specified size from the target population. We proposed a bootstrap-based algorithm for calculating EVSI and demonstrated its feasibility through a case study. We further examined the relationship between EVSI and future development sample sizes, suggesting that scaling EVSI to the population level can quantify the total increase in clinical utility in terms of true or false positives.
Contemporary approaches for the assessment of uncertainty in the predictions of a model are rooted in classical inferential statistics. The incorporation of decision-theoretic approaches that quantify the expected gain by procuring further development data can provide a complementary perspective. Specifically, we suggest the computation of EVPI (at relevant thresholds) to quantify the expected utility loss due to uncertainty with the current development sample. If this loss was deemed high, the EVSI curve can be generated to evaluate the expected return-on-investment from a future development sample. The EVSI curve can be juxtaposed to multi-criteria sample size formulas for model development (Riley and others 2024). This might demonstrate, for example, if one of the specific criteria falls in the “diminishing return” part of the EVSI curve, and thus enables the investigator team to have an informed discussion about sample size requirements of a future validation study. As a case in point, we go back to our GUSTO-I case study. There, applying the equations by Riley and others (2024) at the threshold value of $z=0.02$ results in future sample size estimates of 159,170 for the observed/expected (O/E) criterion (targeting SE of 0.051), 2,477 for the C-slope criterion (targeting SE of 0.051), and 30,938 for the C-statistic criterion (targeting SE of 0.026). Notably, two out of three of these estimates, along with the final recommendation, fall within the “diminishing return” region ($n>10,000$), where further increasing the sample size would provide minimal benefit according to EVSI (Sadatsafavi and others, 2023b). Based on this, investigators might be able to justify settling on the sample sizes determined by the other two criteria.
Recently, the impact of prediction uncertainty has been assessed in terms of the instability of predicted risks (Riley and Collins 2023). While prediction instability is indeed a relevant concerned, its impact on the overall utility of subsequent decisions is not clear. As well, to what extent a future development sample of a given size will reduce instability was not studied by the investigators. We note that the bootstrap-based approach for quantifying prediction instability, as proposed by Riley and Collins (2023) et al, is similar in operation to the bootstrap-based algorithms for VoI analysis. As such, one set of bootstrap-based simulations can be used for the evaluation of prediction instability, EVPI, and EVSI.
This research represents a continued effort to address prediction model uncertainty due to finite sample sizes (Sadatsafavi and others 2022; Sadatsafavi and others 2023a; Sadatsafavi and others 2023b). However, prediction models are also subject to other sources of uncertainty, such as measurement errors in predictors and outcomes or potential discrepancies between the development sample and the target population. Future research could explore the impact of other sources of uncertainty (e.g., whether the sample is truly representative of the population, whether future models will include additional covariates). Additionally, since EVSI itself is derived from a finite sample, it is subject to similar uncertainties as the prediction model. Assessing EVSI uncertainty, potentially by incorporating an additional level of bootstrapping into our EVSI proposed algorithm, could be another avenue for future work. Moreover, the present method assumes the availability of pilot data from the target population, and future studies should investigate how EVSI methods could be developed for broader contexts, including situations where no pilot data are available. This would enhance the integration of EVSI with sample size calculations (e.g., Riley and others 2024), potentially providing more flexible and generalizable approaches. Feature selection is another critical step in prediction model development, with various methods, such as LASSO, being used to refine models by regularizing and eliminating parameters based on a tuning parameter. Further research is needed to assess how model development processes, like feature selection, influence EVSI calculations. More importantly, EVSI could be incorporated into feature selection (generating EVSI values fro models of different complexity). This will help model developers in their quest to strive a balance between model accuracy and parsimony, guiding investigators to focus on collecting only the most relevant predictors in future model development. By doing so, EVSI can help reduce costs associated with prediction model development by tailoring both the number of observations and the number of predictors in future samples.
In conclusion, VoI analysis during model development can inform the investigators to evaluate the adequacy of the sample at hand (EVPI) and the desired sample size of a future model development (EVSI). In general, VoI methods can be a promising addition to the entire lifecycle of prediction models, from development to validation and implementation, as they augment classical methods of sample size calculations with decision theory
Acknowledgments
Conflict of Interest: None declared.